Given the total radiant exitance of 64 W∙m-2 integrated over the wavelength range 2.41-2.90 μm, what is the average radiant exitance over this interval in W∙m-2∙μm-1? 


1
Expert's answer
2021-10-12T12:33:22-0400

Average radiant existance = "\frac{Radiant \; existance}{Difference \;in \;Wavelength}"

"= \frac{64 \;W \cdot m^{-2}}{(2.90-2.41) \; \mu m} \\

= \frac{64\;W \cdot m^{-2}}{0.49 \; \mu m} \\

= 130.6122 \; W \cdot m^{-2} \cdot \mu m^{-1}"